OpenStudy (anonymous):

Solve 3x^2 = –12x – 15

OpenStudy (anonymous):

(3x -3)(x+5) = 0 x = 1 or x = -5

OpenStudy (anonymous):

careful, your sign on the 15 turns out to be incorrect then

OpenStudy (anonymous):

try this instead... (see attached)

OpenStudy (anonymous):

sorry that wrong

OpenStudy (anonymous):

3x^2 = –12x – 15 so 3x^2 + 12x + 15 = 0 this has no real roots - have u learned about complex numbers?

OpenStudy (anonymous):

oser's solution is correct, it has complex roots -2+i and -2 -i